OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (−1, 3) and (1, 7)?

OpenStudy (anonymous):

The slope (k) is rise over run, so k = (7-3)/(1-(-1)) = 4/2=2 We also know that the equation of a line is y = kx + m If we plug in a point on the line (1,7) we get: 7 = k+m and since we know what k is we get that m = 7-k = 7-2 = 5 Therefore the equation of that line is: y = kx + m = 2x+5

OpenStudy (anonymous):

in point slope form

OpenStudy (anonymous):

my choices are y − 7 = 4(x − 1) y − 7 = 2(x − 1) y − 3 = 2(x − 1) y − 3 = 4(x + 1)

OpenStudy (anonymous):

@sbsin

OpenStudy (anonymous):

Ok, so point-slope form is generally y – y1 = m(x – x1) we have calculated m which I called k so m = 2, then you can choose the point (x1,y1)=(1, 7): So its y - 7 = 2 (x - 1)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

If you want there's a great video going in to more details here: https://www.youtube.com/watch?v=eHPTyYbNmx4